Teachers’ Retirement System of the City of New York offers several types of investments for its members. Among the choices are investments with fixed and variable rates of return. There are several categories of variable-return investments. The Diversified Equity Fund consists of investments that are primarily made in stocks, and the Stable-Value Fund consists of investments in corporate bonds and other types of lower-risk instruments. The data in TRSNYC represent the value of a unit of each type of variable-return investment at the beginning of each year from 1984 to 2015. (Data extracted from “Historical Data-Unit Values, Teachers’ Retirement System of the City of New York,” bit.ly/SESJF5.)
a.     For each time series, plot the data in a line graph.
b.     For the Diversified Fund, compute a linear trend forecasting equation.
c.     For the Diversified Fund, compute an exponential trend forecasting equation.
d.     For each of the models in parts b and c, forecast the unit value for 2016
e.     Which of the two forecasts do you believe is more accurate and why?
Year Diversified Equity Stable-Value
1984 13.111 10.342
1985 13.176 11.073
1986 16.526 11.925
1987 18.652 12.694
1988 15.564 13.352
1989 20.827 13.919
1990 24.738 14.557
1991 22.678 15.213
1992 28.549 15.883
1993 29.829 16.510
1994 32.199 16.970
1995 30.830 17.351
1996 39.644 17.682
1997 45.389 18.004
1998 54.882 18.341
1999 64.790 18.678
2000 74.220 18.962
2001 67.534 19.320
2002 57.709 19.673
2003 44.843 19.735
2004 55.993 19.609
2005 60.909 19.520
2006 63.038 19.452
2007 70.413 19.441
2008 71.460 19.493
2009 42.493 19.496
2010 52.629 19.374
2011 58.207 19.223
2012 55.630 19.470
2013 62.174 19.382
2014 77.656 19.039
2015 81.363 18.311
Answer #1

In the picture given below in the table in excel sheet t denotes year and Y denotes Diversified equity.

-406.4361 -382.1841 446.2729 466.3625 358523 437.4 441 470 361 -385.5 285 428.2 225 -387.8 169 -354.2 121 277.581 277.5 49 22Diversified Equity 90 70 60 50 30 20 10 o OWe shift th e onigio iz. 199g and 2 2000 Let the lin ec equity) and a be i Σχ so. a and b are estimated by 32 -10705. 28-0. 9d) forcasting value 2016 Lineon trend equalions we bave yt- 4S. 864 2 2 + 0.981056 2e wben t= 2016 i-we-get-from-a χ= 2(2016-